OpenStudy (maheshmeghwal9):

When we divide 50 in two parts , the sum of their reciprocals is 1/12. Find the two parts.

OpenStudy (experimentx):

\[ \frac{1}{x} + \frac{1}{50-x} = \frac{1}{42}\]

OpenStudy (maheshmeghwal9):

not 42 but 12

OpenStudy (experimentx):

sorry ... that's 12 ... not 42

OpenStudy (maheshmeghwal9):

will u find x please coz i have solved this but not getting the answer

OpenStudy (maheshmeghwal9):

what do u get?

OpenStudy (experimentx):

it's a quadratic equation http://www.wolframalpha.com/input/?i=solve+1%2Fx+%2B+1%2F%2850-x%29+%3D+1%2F12

OpenStudy (maheshmeghwal9):

ya ok but answer is given 2/3

OpenStudy (maheshmeghwal9):

i think answer is wrong

OpenStudy (maheshmeghwal9):

u r right

OpenStudy (experimentx):

the two parts are x:(50-x) = 2:3 or 3:2

OpenStudy (maheshmeghwal9):

oh ya i see! what a mind

OpenStudy (ajprincess):